# Sassha: Sharpness-aware Adaptive Second-order Optimization with Stable Hessian Approximation

### v1

@@ Appendix B Algorithm Comparison
Our method Sassha adds perturbation $\epsilon^{\star}_{t}$ before computing the gradient and Hessian estimation to penalize sharpness during the training process for improved generalization–an approach not previously explored in the literature. For stability, we additionally introduce two techniques: an absolute function and a square root to the Hessian estimates. The absolute function enforces Hessian estimates to be semi-positive definite while preserving their magnitude. Also, the square root smoothly adjusts underestimated curvature, stabilizing the Hessian estimates. Consequently, the blend of sharpness reduction and Hessian stabilization enables the efficient reuse of previously computed Hessians, resulting in a stable and efficient algorithm.
@@ Appendix C Convergence Analysis of Sassha
where $\operatorname{diag}$ extracts the diagonal elements of a matrix as a vector, or constructs a diagonal matrix from a vector, and $\epsilon$ is a damping constant. Define $h_{t}$ as
@@
Appendix D Experiment Setting
@@ D.3 Label Noise
Appendix E More Ablations
E.1 Square Root Function
Table 14: Comparison of square-root against damping and clipping.
@@ E.1 Square Root Function
We conduct an ablation study to support our use of the square-rooted preconditioner in Sassha, comparing it to other alternatives to stabilize the preconditioner such as damping or clipping. We search damping and clipping hyperparameters over $\{10^{-4},10^{-6},10^{-8},10^{-12}\}$ and $\{0.1,0.05,0.01,0.005,0.001,0.0005,0.0001\}$, respectively. We note that the square-root employed in Sassha does not require such extensive hyperparamter search. The results are presented in Table 14.
Our experiments demonstrate that the square-rooted preconditioner achieves higher validation accuracy than those with damping or clipping, even with a three times smaller hyperparameter search budget. We provide two possible explanations for this observation. First, clipping and damping rigidly transform Hessian estimates. Specifically, damping shifts the Hessian estimate by a fixed damping factor, and clipping replaces certain Hessian entries with a predefined constant. Without careful tuning, these inflexible modification may lead to incorrect updates in specific directions, degrading performance. In contrast, the square root smoothly adjusts Hessian estimates, preserving their structural integrity while mitigating extreme values. Second, square-rooted preconditioner can be interpreted as the result of a geometric interpolation between the identity matrix $I$ and $H^{\alpha}$. This interpolation has been demonstrated to enable selecting an optimal preconditioner that balances the bias and the variance of the population risk, thereby minimizing generalization error (Amari et al., 2021). In general, $\alpha=1/2$ (i.e., square root) has consistently shown moderate performance across various scenarios (Amari et al., 2021; Duchi et al., 2011; Kingma & Ba, 2015).
E.2 Absolute Value Function
@@
Appendix F Validation Loss Curve for Vision Task
@@
Appendix G Comparison with First-order Baselines with Given More Training Budget than Sassha
@@
Appendix H Effectiveness of Stable Hessian Approximations in Sassha
@@
We observe that Sophia-H (with SAM) performs worse than SAM, whereas Sassha outperforms both methods, validating the effectiveness of the design choices made in Sassha. We attribute this to the reduced compatibility of Sophia-H with SAM compared to Sassha. First, when curvature is reduced due to SAM, clipping may cause a significant increase in the number of Hessian entries replaced by a very small constant. This situation raises the sensitivity to hyperparameters like the clipping threshold and makes the optimization process more dependent on careful tuning. Conversely, the stable Hessian approximation in Sassha, incorporating the absolute function and square rooting, avoids enforcing hard adjustments to the magnitude or direction of the update vector or to Hessian entries. Instead, it smoothly adjusts Hessian, preserving its structural integrity while mitigating extreme values.
In addition, the use of clipping results in Sophia partially performing signSGD over a subset of parameters (Liu et al., 2024), which may lead to suboptimal convergence in typical situations (Karimireddy et al., 2019).
Appendix I Comparison with Advanced SAM Variants
@@
Appendix J Additional Label Noise Experiments
@@
Appendix K M-Sassha: Efficient Perturbation
@@
which entirely eliminates the need for additional gradient computation with similar generalization improvement as the original SAM. We call this low-computation alternative as M-Sassha and evaluate this across vision, language, and label noise tasks, as we did in the main sections. The results are presented in Tables 21, 22 and 23, respectively.

### v2

@@ Appendix B Algorithm Comparison
Our method, Sassha, adds a perturbation $\epsilon^{\star}_{t}$ before computing the gradient and diagonal Hessian estimation to penalize sharpness during training for promoting improved generalization – an approach that, to our knowledge, has not been previously explored in the literature. However, naive second-order optimization under sharpness minimization can be numerically unstable, because, when the overall curvature is small due to sharpness reduction, Hessian underestimation can cause the optimization to yield extremely large update. To address this issue, we introduce two simple techniques to Hessian estimates: the square root and the absolute function. The square root smoothly adjusts underestimated curvature while preserving the relative scale among Hessian entries. The absolute function enforces Hessian estimates to be semi-positive definite while maintaining their magnitude. This not only prevents convergence to saddle points or local maxima, but also allows the square root to operate on Hessian entries retaining the original Hessian magnitude. Together, the combination of sharpness minimization and Hessian stabilization enables efficient reuse of previously computed Hessians, resulting in a stable and efficient algorithm.
@@ Appendix C Convergence Analysis of Sassha
where $\operatorname{diag}$ extracts the diagonal elements of a matrix as a vector, or constructs a diagonal matrix from a vector, and $\epsilon$ is a small constant introduced to prevent division by zero. Define $h_{t}$ as
@@
Appendix D Linear stability analysis on Sassha
In this section, we provide the detailed proof of Section 4.6.
We begin by considering the minimization of the training error
$f(x)=\frac{1}{n}\sum^{n}_{i=1}f_{i}(x)$
via a general second-order optimization method:
$x_{t+1}=x_{t}-P(x_{t};\xi_{t})^{-1}\nabla f(x_{t};\xi_{t})$ (12)
where $\xi_{t}$ is a i.i.d. random variable independent of $x_{t}$. For Sassha, the update is specified as
$x_{t+1}=x_{t}-\eta\,\frac{1}{\sqrt{|\operatorname{diag}\bigl{(}\nabla^{2}f_{\xi_{t}}(x_{t}+\rho\,\nabla f_{\xi_{t}}(x_{t})\bigr{)}|}+\epsilon}\odot\nabla f_{\xi_{t}}(x_{t}+\rho\,\nabla f_{\xi_{t}}(x_{t})),$ (13)
where $\operatorname{diag}$ extracts the diagonal elements of a matrix as a vector, or constructs a diagonal matrix from a vector, and $\epsilon$ is a small constant introduced to prevent division by zero.
We now derive the linear stability conditions for the fixed points of Sassha.
Definition D.1.
(Fixed point). A point $x^{\star}$ is called a fixed point of the stochastic dynamics (12) if, for any $\xi$, we have $\nabla f_{\xi}(x^{\star})=0$.
Definition D.2.
(Linearized Sassha). Let $x^{\star}$ be the fixed point of interest and assume $f(x^{\star})=0$. Consider the quadratic approximation of $f$ near $x^{\star}$: $f_{\xi}(x)\approx\frac{1}{2}(x-x^{\star})^{\top}H_{\xi}(x-x^{\star})$ with $H_{\xi}=\nabla^{2}f_{\xi_{t}}(x^{\star})$. The corresponding linearized Sassha is given by
$x_{t+1}=x_{t}-\eta\frac{1}{\sqrt{\operatorname{diag}(H_{\xi_{t}})}+\epsilon}\odot H_{\xi_{t}}(\tilde{x}_{t}-x^{\star})$ (14)
where $\tilde{x}_{t}=x_{t}+\rho H_{\xi_{t}}(x_{t}-x^{\star})$ is the linearized perturbed point.
For brevity, we define
$d_{\xi_{t}}:=\sqrt{\operatorname{diag}\bigl{(}H_{\xi_{t}}\bigr{)}}+\epsilon,$
Then, the update (14) can be rewritten as
$x_{t+1}=x_{t}-\eta\frac{1}{d_{\xi_{t}}}\odot H_{\xi_{t}}(\tilde{x}_{t}-x^{\star}),$ (15)
Remark D.3.
In a neighborhood of $x^{\star}$, the inverse scaling vector is uniformly upper bounded by
$\frac{1}{d_{\xi_{t}}}\preceq\frac{1}{\epsilon}\mathbf{1}$ (16)
where $\mathbf{1}$ is the vector of all ones.
Definition D.4.
(Linear stability). Consider a fixed point $x^{\star}$ of linearized stochastic dynamic such as (14). We say that $x^{\star}$ is linearly stable if there exists a constant $C$ such that
$\mathbb{E}\bigl{[}\lVert x_{t}-x^{\star}\rVert^{2}\bigr{]}\leq C\lVert x_{0}-x^{\star}\rVert^{2},\text{ for all }t>0$
Theorem D.5.
Assume without loss of generality that $x^{\star}=0$. Then, the fixed point $x^{\star}$ of Sassha is linearly stable if:
$\lambda_{max}\Bigl{(}\bigl{(}I-\frac{\eta}{\epsilon}H-\frac{\eta\rho}{\epsilon}H^{2}\bigr{)}^{2}+\frac{\eta^{2}-2\eta\rho\epsilon}{\epsilon^{2}}\,\bigl{(}\mathbb{E}H_{\xi_{t}}^{2}-H^{2}\bigr{)}+\frac{2\eta^{2}\rho}{\epsilon^{2}}\,\bigl{(}\mathbb{E}H_{\xi_{t}}^{3}-H^{3}\bigr{)}+\frac{\eta^{2}\rho^{2}}{\epsilon^{2}}\,\bigl{(}\mathbb{E}H_{\xi_{t}}^{4}-H^{4}\bigr{)}\Bigr{)}\leq 1.$
Proof.
Our goal is to obtain a bound of the form $\mathbb{E}\lVert x_{t}\rVert^{2}\leq C\lVert x_{0}\rVert^{2}$, for some constant $C$. We begin by substituting (15) into $\mathbb{E}\lVert x_{t+1}\rVert^{2}$ and proceed to expand the terms as follows:
$\mathbb{E}\Bigl{[}\lVert x_{t+1}\rVert^{2}\Bigr{|}x_{t}\Bigr{]}$ $=x_{t}^{\top}\mathbb{E}\Bigl{[}\bigl{(}I-\eta\operatorname{diag}(\frac{1}{d_{\xi_{t}}})H_{\xi_{t}}(I+\rho\,H_{\xi_{t}})\bigr{)}^{\top}\bigl{(}I-\eta\operatorname{diag}(\frac{1}{d_{\xi_{t}}})H_{\xi_{t}}(I+\rho\,H_{\xi_{t}})\bigr{)}|x_{t}\Bigr{]}\,x_{t}$
[…]
$=x_{t}^{\top}\mathbb{E}\Bigl{[}I-\frac{2\eta}{\epsilon}H_{\xi_{t}}(I+\rho\,H_{\xi_{t}})+\frac{\eta^{2}}{\epsilon^{2}}H_{\xi_{t}}^{2}(I+\rho\,H_{\xi_{t}})^{2}|x_{t}\Bigr{]}\,x_{t}$
$=x_{t}^{\top}\mathbb{E}\Bigl{[}I-\frac{2\eta}{\epsilon}H_{\xi_{t}}-\frac{2\eta\rho}{\epsilon}\,H_{\xi_{t}}^{2}+\frac{\eta^{2}}{\epsilon^{2}}H_{\xi_{t}}^{2}+\frac{2\eta^{2}\rho}{\epsilon^{2}}H_{\xi_{t}}^{3}+\frac{\eta^{2}\rho^{2}}{\epsilon^{2}}H_{\xi_{t}}^{4}|x_{t}\Bigr{]}\,x_{t}$
$=x_{t}^{\top}\Bigl{(}I-\frac{2\eta}{\epsilon}H+\frac{\eta^{2}-2\eta\rho\epsilon}{\epsilon^{2}}\,H^{2}+\frac{2\eta^{2}\rho}{\epsilon^{2}}\,H^{3}+\frac{\eta^{2}\rho^{2}}{\epsilon^{2}}\,H^{4}$
$+\frac{\eta^{2}-2\eta\rho\epsilon}{\epsilon^{2}}\,\bigl{(}\mathbb{E}H_{\xi_{t}}^{2}-H^{2}\bigr{)}+\frac{2\eta^{2}\rho}{\epsilon^{2}}\,\bigl{(}\mathbb{E}H_{\xi_{t}}^{3}-H^{3}\bigr{)}+\frac{\eta^{2}\rho^{2}}{\epsilon^{2}}\,\bigl{(}\mathbb{E}H_{\xi_{t}}^{4}-H^{4}\bigr{)}\Bigr{)}\,x_{t}$
$=x_{t}^{\top}\Bigl{(}\bigl{(}I-\frac{\eta}{\epsilon}H-\frac{\eta\rho}{\epsilon}H^{2}\bigr{)}^{2}+\frac{\eta^{2}-2\eta\rho\epsilon}{\epsilon^{2}}\,\bigl{(}\mathbb{E}H_{\xi_{t}}^{2}-H^{2}\bigr{)}+\frac{2\eta^{2}\rho}{\epsilon^{2}}\,\bigl{(}\mathbb{E}H_{\xi_{t}}^{3}-H^{3}\bigr{)}+\frac{\eta^{2}\rho^{2}}{\epsilon^{2}}\,\bigl{(}\mathbb{E}H_{\xi_{t}}^{4}-H^{4}\bigr{)}\Bigr{)}\,x_{t}$
Since for any $x$ and any matrix $A$ the inequality $x^{\top}Ax\leq\lambda_{max}(A)\lVert x\rVert^{2}$ with $\lambda_{max}(A)$ denoting the maximum eigenvalue of $A$, holds true, applying this inequality and taking the total expectation yields the following:
$\mathbb{E}\lVert x_{t+1}\rVert^{2}\leq\lambda_{max}\Bigl{(}\bigl{(}I-\frac{\eta}{\epsilon}H-\frac{\eta\rho}{\epsilon}H^{2}\bigr{)}^{2}+\frac{\eta^{2}-2\eta\rho\epsilon}{\epsilon^{2}}\,\bigl{(}\mathbb{E}H_{\xi_{t}}^{2}-H^{2}\bigr{)}+\frac{2\eta^{2}\rho}{\epsilon^{2}}\,\bigl{(}\mathbb{E}H_{\xi_{t}}^{3}-H^{3}\bigr{)}+\frac{\eta^{2}\rho^{2}}{\epsilon^{2}}\,\bigl{(}\mathbb{E}H_{\xi_{t}}^{4}-H^{4}\bigr{)}\Bigr{)}\,\mathbb{E}\lVert x_{t}\rVert^{2}$
Recursively applying this bound gives
$\mathbb{E}\lVert x_{t}\rVert^{2}=\lambda_{max}\Bigl{(}\bigl{(}I-\frac{\eta}{\epsilon}H-\frac{\eta\rho}{\epsilon}H^{2}\bigr{)}^{2}+\frac{\eta^{2}-2\eta\rho\epsilon}{\epsilon^{2}}\,\bigl{(}\mathbb{E}H_{\xi_{t}}^{2}-H^{2}\bigr{)}+\frac{2\eta^{2}\rho}{\epsilon^{2}}\,\bigl{(}\mathbb{E}H_{\xi_{t}}^{3}-H^{3}\bigr{)}+\frac{\eta^{2}\rho^{2}}{\epsilon^{2}}\,\bigl{(}\mathbb{E}H_{\xi_{t}}^{4}-H^{4}\bigr{)}\Bigr{)}^{t}\,\mathbb{E}\lVert x_{0}\rVert^{2}$
Here, we can see that $x^{*}$ is linearly stable if
$\lambda_{max}\Bigl{(}\bigl{(}I-\frac{\eta}{\epsilon}H-\frac{\eta\rho}{\epsilon}H^{2}\bigr{)}^{2}+\frac{\eta^{2}-2\eta\rho\epsilon}{\epsilon^{2}}\,\bigl{(}\mathbb{E}H_{\xi_{t}}^{2}-H^{2}\bigr{)}+\frac{2\eta^{2}\rho}{\epsilon^{2}}\,\bigl{(}\mathbb{E}H_{\xi_{t}}^{3}-H^{3}\bigr{)}+\frac{\eta^{2}\rho^{2}}{\epsilon^{2}}\,\bigl{(}\mathbb{E}H_{\xi_{t}}^{4}-H^{4}\bigr{)}\Bigr{)}\leq 1.$
Appendix E Experiment Setting
@@ E.3 Label Noise
Appendix F More Ablations
F.1 Square Root Function
Table 14: Comparison of square root against damping and clipping.
@@ F.1 Square Root Function
We conduct an ablation study to support our use of the square rooted preconditioner in Sassha, comparing it to other alternatives to stabilize the preconditioner such as damping or clipping. We search damping and clipping hyperparameters over $\{10^{-4},10^{-6},10^{-8},10^{-12}\}$ and $\{0.1,0.05,0.01,0.005,0.001,0.0005,0.0001\}$, respectively. We note that the square-root employed in Sassha does not require such extensive hyperparamter search. The results are presented in Table 14.
[…]
F.2 Absolute Value Function
@@
Appendix G Validation Loss Curve for Vision Task
@@
Appendix H Comparison with First-order Baselines with Given More Training Budget than Sassha
@@
Appendix I Effectiveness of Stable Hessian Approximations in Sassha
@@
We observe that Sophia-H (with SAM) performs worse than SAM, whereas Sassha outperforms both methods, validating the effectiveness of the design choices made in Sassha. We attribute this to the reduced compatibility of Sophia-H with SAM compared to Sassha. First, Sophia clipping destroys the relative scale between individual elements of the Hessian. This may be particularly problematic when using SAM, where the overall Hessian values tend to be small. In such cases, even very small differences between Hessian elements may carry nontrivial curvature information. However, Sophia clipping abruptly replaces small or negative Hessian values based on a predefined and fixed threshold criterion. As a result, when the Hessian is generally small due to SAM, informative dimensions may fall below the threshold, removing potentially critical variations and thereby deteriorating the quality of preconditioning. This situation also raises the sensitivity to hyperparameters like the clipping threshold and makes the optimization process more dependent on careful tuning. Conversely, the stable Hessian approximation in Sassha, incorporating the absolute function and square rooting, preserves the relative scale among the Hessian entries by smoothly adjusting their magnitudes.
In addition, the use of sophia clipping results in Sophia partially performing signSGD over a subset of parameters (Liu et al., 2024), which may lead to suboptimal convergence in typical situations (Karimireddy et al., 2019).
Appendix J Additional Language Modeling Experiments
Table 17: GPT2 pretraining results. Sassha achieves similar performance to state-of-the-art methods.
Method Loss Perplexity
AdamW 2.9622 19.353
Sophia-G 2.9307 18.751
SAM ${}_{\text{AdamW}}$ 2.9558 19.196
Sophia-G (with SAM) 2.9319 18.773
Sassha 2.9445 19.015
In this section, we provide additional language modeling experiments. We train GPT2-small for 50k iterations on OpenWebText (Gao et al., 2020), using various optimization methods for comparison. For AdamW and Sophia-G, we adopt the best hyperparameter configurations reported by (Liu et al., 2024), and set $\rho=0.1$ for all SAM-related methods. Due to limited computational resources, all experiments were run with a single random seed. The results are presented in Table 17. We observe that Sassha achieves performance comparable to related methods.
Appendix K Comparison with Advanced SAM Variants
@@
Appendix L Additional Label Noise Experiments
@@
Appendix M M-Sassha: Efficient Perturbation
@@
which entirely eliminates the need for additional gradient computation with similar generalization improvement as the original SAM. We call this low-computation alternative as M-Sassha and evaluate this across vision, language, and label noise tasks, as we did in the main sections. The results are presented in Tables 22, 23 and 24, respectively.